# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ 4.1 Governance for Agentic LLM Systems
AI Trust, Risk, and Security Management (AI TRISM) is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards habbal2024artificial . Originally highlighted in industry guidelines for AI governance litan2024ai , TRISM provides a structured approach to manage the unique challenges of LLM-based “agentic” AI systems. Such systems consist of autonomous LLM agents that can make independent decisions, collaborate with other agents, and adapt their behavior over time kaur2024building . These properties: autonomy, multi-agent interaction, and evolving behavior, introduce new risks not seen in traditional single-model deployments ko2025sevensecuritychallengessolved .
For example, an agent acting in isolation might be benign, but when interacting with others across organizational or trust boundaries, it could manipulate peers or leak confidential information. The TRISM framework addresses these concerns by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy kaur2024building ; habbal2024artificial . Each pillar targets a critical aspect of safety or risk management, ensuring that an agentic LLM system remains transparent, reliable, secure, and compliant with ethical and regulatory requirements. Below, we define each pillar and explain how it applies to LLM-based agentic systems, grounding the discussion in current research and best practices.
@@ 4.2 Explainability in Multi-Agent Decision Making
Definition and Importance Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of agentic LLM systems, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction rosenfeld2019explainability . Accordingly, the TRISM framework elevates explainability as a core pillar to ensure that each agent’s actions and the overall system’s behavior can be understood and audited.
Techniques for Explainability in Agentic Systems Achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes. Established explainable AI techniques provide a starting point. For instance, Local Interpretable Model-Agnostic Explanations (LIME) ribeiro2016should and Shapley Additive Explanations (SHAP) lundberg2017unified can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an agentic setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered kusner2017counterfactual , for example, if a particular agent’s contribution were removed or modified. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making. For instance, by systematically toggling an agent off or varying its outputs, one can observe changes in the collective outcome and thus determine that Agent X was critical in influencing decision Y. Such analyses surface the dependencies and influences among agents, effectively explaining emergent behaviors at the system level. Moreover, recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g. chain-of-thought prompts or dialogue between agents) can further enhance transparency. Human auditors can also help with a trace of how agents arrived at a decision, e..g, which agent contributed what information and why, to produce narrative explanations for its outcomes.
Summary In summary, the explainability pillar of TRISM compels the use of these techniques (surrogate models, feature attributions, counterfactual testing, and transparent reasoning traces) to ensure that even highly autonomous LLM agents remain interpretable and accountable to human oversight.
@@ 4.3 ModelOps : Lifecycle Management for LLM Agents
Definition and Scope ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement lefevre2022modelops . It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production. Within agentic LLM systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely kaur2024building . Effective ModelOps is crucial for maintaining consistency, performance, and regulatory compliance as LLM agents evolve or as new agents are added to a system.
Application to Agentic Systems LLM-based agents require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental drift. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions. This aligns with best practices in LLM operations (LLMOps) mlops , which integrate MLOps principles tailored to the challenges of large language models
Challenge One challenge with this method is model drift, i.e., over time, an agent’s responses may become less accurate or relevant as data distributions shift or real-world conditions change. Continuous monitoring is therefore required to detect performance degradation or deviations from expected behavior, triggering retraining or recalibration when needed. Moreover, real-time monitoring and logging are fundamental in multi-agent settings. Each agent’s actions (e.g. API calls, decisions taken, errors encountered) are logged and analyzed to provide observability into the system’s functioning.
In a large-scale agent ecosystem, orchestration services may oversee the agents, scheduling their tasks and managing inter-agent communication. ModelOps must govern these orchestration layers as well, ensuring, for instance, that if one agent fails or produces suspicious output, it can be isolated or shut down without collapsing the whole system.
Summary In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable. It formalizes change management (so updates do not introduce new risks), provides continuous validation of model behavior, and supports compliance by logging data for audits and enforcing policies (e.g. preventing unauthorized model changes).
@@ 4.4 Application Security for LLM Agents
Threats in Agentic Environments Application security in the TRISM framework focuses on protecting AI agents and their ecosystem from malicious attacks and misuse. LLM-based agents are susceptible to a range of novel security threats that exploit their language-based interfaces and cooperative behaviors. One well-documented threat is prompt injection kosinski2024prompt , wherein an attacker designs input data that contains hidden or malicious instructions . Recent studies lee2024promptinfectionllmtollmprompt have shown that in AMAS, such prompt injections can propagate from one agent to another, a phenomenon dubbed “prompt infection” is analogous to a computer virus spreading across networks raza2025justhumansneedvaccines . In a prompt infection scenario, a malicious prompt introduced to Agent A might covertly modify Agent A’s output, which then becomes part of Agent B’s input, thereby tricking Agent B as well, and so on. This cascading attack can lead to widespread data leaks, fraudulent transactions, misinformation, or coordinated misbehavior across an agent society.
Another critical vulnerability is identityspoofing and impersonation chen2025aiagents . In a multi-agent system, agents often communicate or coordinate tasks, and they may rely on credentials or tokens to authenticate one another. For example, if an adversary steals an agent’s API key or tricks the system into treating a rogue model as a trusted peer, they could issue commands or receive information under a false identity.
Defensive Measures: To mitigate these threats, TRISM’s security pillar mandates a defense-in-depth approach tailored to LLM agents:
  1. 1.
    Prompt hygiene: agents should treat inputs defensively by sanitizing and filtering prompts, and using guardrails or content policies to detect and refuse suspicious instructions. Prompt hardening (e.g., adding secure prefixes or validation steps) is one such method to make agents less susceptible to injection gojsalic2024system .
  2. 2.
    Strong authentication and access controls Each agent and human user must be securely authenticated, and least-privilege principles owasp2025agentic should constrain what actions an agent can perform autonomously.
  3. 3.
    Continuous monitoring If an agent suddenly starts issuing unusual requests or deviates from its normal pattern of behavior, automated monitors can flag this for investigation or trigger an automatic shutdown of the agent’s actions.
Recent frameworks, such as LangChain/LangFlow langchain2024function , AutoGen autogen_agents_tutorial , CrewAI crewAI , introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) raza2025justhumansneedvaccines can improve their resistance to malicious inputs
Summary In sum, the Application Security pillar of TRISM emphasizes proactive safeguards against both external attackers and potential rogue agents. By implementing strict authentication, input validation, encrypted communication, sandboxing of execution (for agents that can use tools or code), and comprehensive monitoring, organizations can significantly reduce the risk of prompt-based exploits, impersonation, and other lateral vulnerabilities that are unique to autonomous multi-agent AI systems hal.science. This layered security approach is essential to maintain the integrity and reliability of agentic LLM deployments in adversarial environments.
@@ 4.5 Model Privacy and Data Protection
Challenges in Data Sharing: The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms. LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks ko2025sevensecuritychallengessolved . In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. Therefore, TRISM’s privacy pillar compels organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
Techniques for Privacy Preservation
  • Differential Privacy (DP): Injects calibrated noise during model training to prevent memorization of individual data entries, ensuring no single record significantly influences the output dwork2006calibrating .
  • Data Anonymization and Minimization: Limits inter-agent data sharing to only what is necessary sweeney2002k , often using aggregated or pseudonymized formats (e.g., “age bracket 30-40” instead of exact birthdate).
  • Secure Multi-Party Computation (SMPC): Enables agents to compute joint functions without exposing private inputs, useful in cross-organization tasks like collaborative fraud detection lindell2009secure .
  • Homomorphic Encryption (HE): Allows agents to compute on encrypted data gentry2009fully . With Fully Homomorphic Encryption (FHE), even the plaintext query and response remain unseen by the agent.
  • Trusted Execution Environments (Secure Enclaves): Hardware-based isolation ensures that even privileged system users cannot access the data being processed by agents costan2016intel . Useful for secure memory sharing and execution.
  • Model Privacy Policies and Compliance: Enforces data retention limits, maintains audit logs, and ensures compliance with regulations (e.g., GDPR, HIPAA) governing agent behavior and data usage Responsi73:online .
Summary By implementing these layers of privacy defense , from differential privacy in model training to homomorphic encryption for data sharing, and stringent access control policies, Agentic AI systems can protect user data and proprietary information even as they leverage that data for intelligent decision-making.
The TRISM framework offers a comprehensive governance model for LLM-based Agentic AI, integrating Explainability, ModelOps, Application Security, and Model Privacy to manage the complexity of autonomous agent systems. Grounded in proven methods, such as SHAP, CI/CD, adversarial defenses, and homomorphic encryption, TRISM enhances safety, transparency, and trust. As AI systems evolve, TRISM provides a stable foundation to ensure responsible and secure agent behavior, aligning advanced capabilities with human values and operational integrity.
@@
Table 4: TRISM pillars for LLM-based Agentic AI and associated governance references.
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” dotan2024evolving
ISO/IEC 24029-1 iso24029-1
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) iso42001
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps owasp2024llmtop10
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 gdpr25
HIPAA §164 hipaa164
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
5 Threats and Risks in LLM-based (Agentic) AMAS
In this section, we discuss the threats and risks in AMAS.
5.1 Unique Threat Vectors
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures deng2025ai . These risks primarily arise from agents’ autonomy, persistent state management, and the complex demands of multi-agent coordination sapkota2025ai . Below, we discuss these threats
  • Autonomy abuse The foremost threat is autonomy abuse, whereby agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs patlan2025real ; kolt2025governing . Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states hammond2025multi ; wilson2025age .
  • Persistent memory Another threat is the persistent memory, which while, crucial for context retention, introduces unique vulnerabilities through potential adversarial injections and accumulations du2025rethinking ; dechant2025episodic . Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms patlan2025real ; deng2025ai .
  • Agent orchestration This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT hong2023metagpt and AutoGen wu2023autogen . These orchestration vulnerabilities differentiate agentic systems markedly from conventional stateless, single-threaded LLM deployments.

### v2

@@ 5 The TRiSM Framework for Agentic AI
The TRiSM Framework for Agentic AI is presented in this section.
Table 4: TRiSM pillars in Agentic AI systems with aligned governance references.
Explainability Transparent multi-agent decisions SHAP, LIME, counterfactuals; reasoning-trace logs; cross-agent dependency mapping NIST AI RMF “Explain” [87]
ISO/IEC 24029-1 [88]
Metrics: fidelity, sufficiency, human-trust score
ModelOps Lifecycle governance of agents, models, orchestration Version-controlled prompts; CI/CD w/ multi-agent sims; drift monitoring; rollback policies ISO/IEC 42001 (AI MS) [89]
SRE MTTR/MTTD; $\Delta$-accuracy (model drift); pipeline pass-rate
Application Security Prevent prompt injection, spoofing, lateral exploits Prompt sanitation; least-privilege scopes; encrypted comms; sandboxed tool calls; anomaly detectors OWASP Top-10 for LLM Apps [90]
Jailbreak-success%; mean-time-to-detect; exploit CVSS
Model Privacy Safeguard sensitive data and shared memory Differential-privacy training; data minimization; homomorphic encryption; secure enclaves; audit logs GDPR Art. 25 [91]
HIPAA §164 [92]
Metrics: $\epsilon$-DP budget, leakage rate, access-audit pass-rate
@@ 5.1 Governance for Agentic LLM Systems
AI TRiSM is a comprehensive governance framework designed to ensure that AI systems are trustworthy, robust, and compliant with safety standards [29]. Originally highlighted in industry guidelines for AI governance [93], TRiSM provides a structured approach to manage the unique challenges of LLM-based “ Agentic AI" systems (herein we refer to as Agentic LLM systems). The TRiSM framework addresses the trust concerns in Agentic AI by focusing on four key pillars: Explainability, Model Operations (ModelOps), Application Security, and Model Privacy [59, 29]. Figure 4 illustrates how these four TRiSM pillars map onto the core architectural layers of an LLM-based Agentic AI system.
Figure 4: TRiSM-Aligned Architecture for LLM-Driven Agentic AI Systems. The system is organised into four vertical layers: Trust & Explainability, ModelOps, Planning + Reasoning, and Model Privacy, each supported by a cross-cutting Monitoring & Governance Plane that enforces policies, logging, and compliance triggers. External tools are accessed through a single Secure API gateway, while the user interface feeds back into both the Trust layer and the governance plane to enable end-to-end oversight.
The TRiSM framework aligns closely with several key international AI governance efforts that shape best practices for safety, accountability, and transparency in AI systems. For example, the EU AI Act [94] (notably Article 14) mandates clear obligations for transparency, traceability, and human oversight in high-risk systems, principles directly reflected in TRiSM Explainability and ModelOps pillars. The NIST AI Risk Management Framework (AI RMF) [95] emphasizes Explainability, Governability, and Validity as core risk mitigation strategies, providing a conceptual match to TRiSM’s structured controls across lifecycle and system interaction points. Additionally, the OECD AI Principles advocate for transparency, robustness, and accountability [96]each echoed in TRiSM attention to secure tool invocation, memory traceability, and multi-agent auditability. These global frameworks reinforce the need for Agentic AI systems to support real-time monitoring, enforce role-based access, maintain verifiable logs, and provide intelligible outputs. As shown in Figure 5, TRiSM pillars provide operational scaffolding for implementing regulatory expectations in LLM-based multi-agent contexts.
Alignment of TRiSM Pillars with Governance Frameworks Explainability: Aligned with the EU AI Act (Art. 13–14) [94] and the NIST AI RMF “Explain” function [95]. This pillar promotes transparent, rationale, human-interpretable logging, and role-based interpretability [97]. ModelOps (Lifecycle Governance): Anchored in the NIST RMF “Govern” function [95] and ISO/IEC 42001 [98]. It emphasizes lifecycle controls, model versioning, rollback policies, and CI/CD pipelines with safety gates. Application Security: Guided by OWASP LLM Top-10 [90] and OECD AI Principles on Robustness [96], it covers access controls, prompt sanitization, sandboxing, and audit trail mechanisms. Model Privacy: Aligned with GDPR (Art. 25) [91], HIPAA §164 [92], and OECD AI Principles [96], this pillar ensures data minimization, encryption, differential privacy budgeting, and access-audit pass thresholds.
Figure 5: Governance alignment of TRiSM pillars with international frameworks such as the EU AI Act, NIST RMF, ISO/IEC 42001, and OECD Principles. Each pillar operationalizes specific compliance strategies to ensure transparency, resilience, and trust in LLM-based systems.
@@ 5.2 Explainability in Multi-Agent Decision Making
Explainability refers to making the inner workings and decisions of AI agents interpretable to humans. In the context of AMAS, explainability is paramount for building user trust, as outcomes often emerge from complex interactions among multiple agents rather than a single model’s prediction [99]. However, achieving explainability in multi-agent LLM systems is challenging because one must interpret not only individual model decisions, but also the inter-agent dynamics that lead to final outcomes.
Established explainable AI techniques provide a starting point. For instance, LIME [100] and Shapley Additive Explanations (SHAP) [101] can be adapted to analyze LLM decisions. These techniques identify which features or input factors most influenced an agent’s output, offering insight into why a particular action or response was taken. In an Agentic LLM setting, a “feature” might be a component of the agent’s input context or a signal from another agent.
Beyond local explanations, counterfactual analysis is increasingly important for multi-agent explainability. Counterfactual techniques examine how the system’s behavior would change if certain conditions were altered [102], for example, if a particular agent’s contribution were removed or modified. This approach, rooted in causal inference, helps isolate each agent’s role in collaborative decision-making. Recent research on explainable AMAS suggests logging the intermediate reasoning steps (e.g. CoT prompts or dialogue between agents) can further enhance transparency. Human auditors can also help with a trace of how agents arrived at a decision, e.g., which agent contributed what information and why, to produce narrative explanations for its outcomes.
@@ 5.3 ModelOps : Lifecycle Management for LLM Agents
ModelOps is the discipline of managing AI models through their entire lifecycle, from development and deployment to monitoring, maintenance, and eventual retirement [103]. It extends the principles of MLOps (Machine Learning Operations) to focus specifically on model governance and reliable operation in production. Within Agentic AI systems, ModelOps encompasses not just individual models, but the orchestration of multiple agents and the supporting infrastructure that keeps them running safely [59]. Effective ModelOps is important to maintain consistency, performance, and regulatory compliance as LLM agents evolve or are added to a system.
Agentic AI systems require rigorous lifecycle governance because their behavior can change with model updates, prompt adjustments, or environmental variations. A cornerstone of ModelOps in this context is version control, i.e., tracking and managing versions of each agent’s model and prompt configurations. Additionally, robust CI/CD pipelines (Continuous Integration/Continuous Deployment) are employed to automatically test agents’ performance and safety whenever a model is fine-tuned or an agent’s logic is modified. Before deployment, multi-agent simulations and unit tests validate that new agent behaviors do not introduce regressions or unsafe interactions. This aligns with best practices in LLM operations (LLMOps) [104], which integrate MLOps principles tailored to the challenges of LLMs. In line with TRISM objectives, robust ModelOps thereby ensures that an Agentic AI system remains reliable and maintainable.
@@ 5.4 Application Security for LLM Agents
To mitigate threats such as prompt injection, spoofing, and impersonation, the TRiSM security pillar recommends a defense-in-depth approach tailored to LLM agents. This includes prompt hygiene [25] , where agents treat inputs defensively by sanitizing and filtering prompts, and applying guardrails or content policies to detect and refuse suspicious instructions. Techniques like prompt hardening, such as secure prefixes or validation steps, can help reduce susceptibility to injection attacks [105]. It also involves strong authentication and access controls, ensuring that each agent and human user is securely authenticated, and that least-privilege principles are enforced to limit autonomous agent actions [106]. Finally, continuous monitoring plays a critical role: if an agent begins issuing anomalous requests or deviates from typical behavior patterns, automated systems can flag the incident for review or trigger a shutdown to prevent further risk.
Recent frameworks, such as LangChain [107], AutoGen [108], and CrewAI [81] introduce the idea of trust scores or reputation among agents, where agents verify each other’s outputs and cross-check decisions to catch inconsistencies or signs of compromise. Moreover, training LLM agents with adversarial robustness in mind (e.g. fine-tuning on adversarial examples, employing adversarial training regimes) [25] can improve their resistance to malicious inputs.
@@ 5.5 Model Privacy and Data Protection
The Model Privacy pillar addresses the protection of sensitive data within AI agent systems, ensuring that the use of personal or confidential information complies with privacy regulations and ethical norms [27]. Since multiple agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate. Without strict privacy controls, there is a risk that an agent could inadvertently expose private data to unauthorized parties or that sensitive information could “leak” through the language model’s outputs. Therefore, TRiSM’s privacy pillar recommends organizations to institute measures that safeguard data throughout the AI lifecycle, from training and inference to inter-agent communication.
To preserve privacy in multi-agent LLM systems, several technical measures are employed. Differential Privacy (DP) adds calibrated noise during model training to ensure that no individual data point can significantly influence the output, thus preventing memorization of personal information [109]. Data anonymization and minimization further enhance privacy by limiting shared information to only what is essential, often in aggregated or pseudonymized formats, such as using “age bracket 30–40” instead of precise birthdates [110]. Secure Multi-Party Computation (SMPC) allows agents to jointly compute functions without revealing their private inputs, making it particularly useful in cross-organizational scenarios like collaborative fraud detection [111]. Similarly, Homomorphic Encryption (HE) permits computation directly on encrypted data; with Fully Homomorphic Encryption (FHE), even plaintext queries and responses remain inaccessible to the agent itself [112]. Trusted Execution Environments (TEEs), or secure enclaves, offer hardware-based isolation, ensuring that even privileged users cannot access data during processing, which is an essential safeguard for secure memory sharing and execution [113]. Finally, the implementation of model privacy policies and compliance frameworks—including data retention limits, audit logs, and adherence to regulations such as GDPR or HIPAA, ensures that agents operate within legal and ethical boundaries [114].
@@
6 Threats and Risks in LLM-based Agentic Multi-Agent Systems (AMAS)
In this section, we present the main threats in Agentic AI and also discuss a taxonomy of emergent risk classes.
Figure 6: Overview of Key Risks and Threats in Agentic AI Systems. The left panel outlines system-level threats unique to agentic architectures, including Autonomy Abuse, Persistent Memory, and Agent Orchestration. The right panel presents a taxonomy of emergent risk classes : Adversarial Attacks, Data Leakage, Agent Collusion, and Emergent Behaviors that arises from multi-agent interactions, memory persistence, and decentralized decision-making.
6.1 Threats in Agentic AI
Agentic AI systems introduce a distinct set of security and reliability concerns compared to traditional single-agent LLM architectures [115]. Below, we discuss these threats.
Autonomy abuse. The foremost threat is autonomy abuse, where agents with significant decision-making authority might misinterpret objectives or implement harmful plans due to erroneous reasoning or manipulated inputs [116]. Unlike deterministic models, agentic systems dynamically generate actions, complicating efforts to define and enforce safe operational states [117].
Persistent memory. Another threat is the persistent memory, which, while crucial for context retention, introduces unique vulnerabilities through potential adversarial injections [118, 119]. Such contamination can propagate subtly via shared memory, especially in the absence of detailed version control and robust audit mechanisms [116].
Agent orchestration. This risk involves central or distributed control mechanisms for role assignment and workflow mediation. A compromised orchestrator could distort task distribution or misroute information, triggering cascading failures, issues exemplified by documented vulnerabilities in MetaGPT [19] and AutoGen [18]. These orchestration vulnerabilities differentiate Agentic AI systems from conventional stateless, single-threaded LLM deployments.